Reports 1-1 of 1 Clear search Modify search
Detector Characterisation (Glitches)
cleva - 13:56 Monday 05 August 2019 (46593) Print this report
Comment to More PSTAB glitches (46564)

# understand why PSL require more corrrection from PStab, understand how PMC intra cavity loss evolve

see plot1 (plot2 gives the legend)

- it is visible that the PMC REFL is worsened at each PSL stops (ENEL on 9/06/19, ML fibering on 30/07/19), see the yellow line at ~ 15 W

- it is visible that PMC REFL is more unstable after the ML fibering, likely due to a residual misalignement (although the fibering should have not affected it)

- it is visible that the PC throughput is worsening since the ENEL shutdown, no idea why. It can be fake and connected to some drifts, since the PMC_temperature (brown red) does not seem to be impacted like it used to be Oct the 28 (see plot 3 & 4, temperature increases by 5 units)

- the extra demand for PSTAB_HF_CORR around the 1/08/18 16H UTC does not seem related to anything specific, looking at the trend of the signals of interest in PSL

- the drift of the noeVAN beam (green curve) continues. We have guessed that this might be connected to the pump diodes ageing. This can induce the noticeable drift in the beam position at the neoVAN output.

# Conclusion

- worth to realign properly the PMC to reduce the overall associated instabilities, might improve the HF_CORR too

- PSL overall alignement needs to be done some time (oct. break). Also to make a direct meas. of the PMC throughput

Images attached to this comment
Search Help
×

Warning

×